OpenStudy (anonymous):

Solve. 3x^2 + 6x + 6 = 0

OpenStudy (turingtest):

first off you can factor out a 3, so try that

Directrix (directrix):

3x^2 + 6x + 6 = 0 3( x^2 + 2x + 2) = 0 x^2 + 2x + 2 = 0 Use the Quadratic Formula to solve for the roots. x = -1 + i and x = -1 + i See the graph at the following link: http://www.wolframalpha.com/input/?i=x%5E2+%2B+2x+%2B+2%3D0